The equilibrium constant for the following reaction is 1.29×10-2 at 600K.

  • COCl2(g)>CO(g) + Cl2(g)


If an equilibrium mixture of the three gases at 600K contains 2.34×10-2 M COCl2(g) and 1.39×10-2 M CO, what is the equilibrium concentration ofCl2? ___M
Answer #1

Kc is defined as concentration of product by concentration of reactant with each concentration term raised to power that is equal to its stoichiometric coefficient in balanced equation

pure liquid and solid do not appear in Kc

So,

Kc = [CO][Cl2]/[COCl2]

1.29*10^-2 = (1.39*10^-2)*[Cl2]/(2.34*10^-2)

[Cl2] = 2.17*10^-2 M

Answer: 2.17*10^-2 M